# Andreas Petrou-Zeniou

### v3

@@ Appendix E Simulation Study Results
In this section, we present the results of a simulation study on our two-step approach to inference on multiple winners. We compare our two-step method to the existing approaches from A, namely the locally simultaneous approach of Zrnic and Fithian (2024b) and test inversion approach of Zrnic and Fithian (2024a). Two-step inference performs well in simulations and relative to these methods, a finding which is robust to correlation.
@@
Figure 7: Coverage probability and length in design 1. CI lengths are presented as fractions of the projection CI length.
Figure 8: Coverage probability and length in design 2. CI lengths are presented as fractions of the projection CI length.
Figure 9: Coverage probability and length in design 3. CI lengths are presented as fractions of the projection CI length.
Figure 10: Coverage probability and length in design 4. CI lengths are presented as fractions of the projection CI length.
Figure 11: Coverage probability and length in design 5. CI lengths are presented as fractions of the projection CI length.
Figure 12: Coverage probability and length in design 6. CI lengths are presented as fractions of the projection CI length.
Figure 13: Coverage probability and length in design 7. CI lengths are presented as fractions of the projection CI length.

### v4

@@ Appendix D Simulation Study Results
In this section, we present the results of a simulation study on our two-step approach to inference on multiple winners. We compare our two-step method to the existing approaches from section 4 and from appendix A, namely the locally simultaneous approach of Zrnic and Fithian (2024b) and test inversion approach of Zrnic and Fithian (2024a). two-step inference performs well in simulations and relative to these methods, a finding which is robust to correlation.
@@
Figure 5: Coverage probability and length in design 1. CI lengths are presented as fractions of the projection CI length. We denote the two-step approach to inference in red, projection in black, the zoom test (based on the step-wise implementation) in light gray, and locally simultaneous inference in dark gray.
Figure 6: Coverage probability and length in design 2. CI lengths are presented as fractions of the projection CI length. We denote the two-step approach to inference in red, projection in black, the zoom test (based on the step-wise implementation) in light gray, and locally simultaneous inference in dark gray.
Figure 7: Coverage probability and length in design 3. CI lengths are presented as fractions of the projection CI length. We denote the two-step approach to inference in red, projection in black, the zoom test (based on the step-wise implementation) in light gray, and locally simultaneous inference in dark gray.
Figure 8: Coverage probability and length in design 4. CI lengths are presented as fractions of the projection CI length. We denote the two-step approach to inference in red, projection in black, the zoom test (based on the step-wise implementation) in light gray, and locally simultaneous inference in dark gray.
Figure 9: Coverage probability and length in design 5. CI lengths are presented as fractions of the projection CI length. We denote the two-step approach to inference in red, projection in black, the zoom test (based on the step-wise implementation) in light gray, and locally simultaneous inference in dark gray.
Figure 10: Coverage probability and length in design 6. CI lengths are presented as fractions of the projection CI length. We denote the two-step approach to inference in red, projection in black, the zoom test (based on the step-wise implementation) in light gray, and locally simultaneous inference in dark gray.
Figure 11: Coverage probability and length in design 7. CI lengths are presented as fractions of the projection CI length. We denote the two-step approach to inference in red, projection in black, the zoom test (based on the step-wise implementation) in light gray, and locally simultaneous inference in dark gray.